OpenStudy (anonymous):

Quick question: Does sec^4(x)-tan^4(x) over sec^2(x)+tan^2(x) simplify into cos^2(x) over cos^2(x), which equals 1?

OpenStudy (anonymous):

WAIT never mind, it's wrong, isn't it?

Directrix (directrix):

I think the 1 is correct. I factored the numerator (difference of two squares) and divided out the denominator of sec² x + tan²x. That left sec² x - tan²x. Then, I used the identity: 1 + tan²x = sec² x to get 1. Check that.

OpenStudy (anonymous):

Could you show me each step you did? I'm still a bit confused

Directrix (directrix):

Factor the numerator: sec^4 (x) - tan^4(x) = [ sec^2(x) + tan^2(x) ] *[ [ sec^2(x) - tan^2(x) ]

OpenStudy (anonymous):

OHHH okay got it. Thank you so much! i understand now.

Directrix (directrix):

Do you see that the denominator of sec^2(x)+tan^2(x) is a factor of the numerator?

OpenStudy (anonymous):

Yes.

Directrix (directrix):

After division, you are left with this: sec^2(x) - tan^2(x)

OpenStudy (anonymous):

Okay.

Directrix (directrix):

Trig Identity: 1 + tan^2(x) = sec^2(x). Learn those identities.

OpenStudy (anonymous):

Alright!

Directrix (directrix):

Take the numerator of sec^2(x) - tan^2(x) and in place of sec^2(x), put 1 + tan^2 (x) to get ....

OpenStudy (anonymous):

sec^2(x), correct?

Directrix (directrix):

1 + tan^2 (x) - tan^2x = 1.

Directrix (directrix):

sec^2(x) - tan^2(x) = (1 + tan^2 (x) ) - tan^2x = 1+ tan^2 (x) - tan^2(x) = 1.

OpenStudy (anonymous):

I understand now. Thank you so much!

Directrix (directrix):

You are welcome.